# Evolving a Multi-Population Evolutionary-QAOA on Distributed QPUs

### v2

@@ Acknowledgment
This work was supported by the Hartree National Centre for Digital Innovation, a UK Government-funded collaboration between STFC and IBM. IBM, the IBM logo, and www.ibm.com are trademarks of International Business Machines Corp., registered in many jurisdictions worldwide. Other product and service names might be trademarks of IBM or other companies. The current list of IBM trademarks is available at www.ibm.com/legal/copytrade.
The research in this paper made use of the following software packages and libraries: Python [34], Qiskit [27], Numpy [35], Scipy [28], Random [36], NetworkX [37, 38], Matplotlib [39, 40].

### v3

@@ Acknowledgment
This work was supported by the Hartree National Centre for Digital Innovation, a UK Government-funded collaboration between STFC and IBM. IBM, the IBM logo, and www.ibm.com are trademarks of International Business Machines Corp., registered in many jurisdictions worldwide. Other product and service names might be trademarks of IBM or other companies. The current list of IBM trademarks is available at www.ibm.com/legal/copytrade. The research in this paper made use of the following software packages and libraries: Python [51], Qiskit [33], Numpy [52], Scipy [34], Random [53], NetworkX [54, 55], Matplotlib [56, 57].